OpenStudy (anonymous):

How would you express this piece-wise function in terms of the unit-step function?

OpenStudy (anonymous):



OpenStudy (anonymous):

I got something like this:\[f\left (t\right )=\mathcal{U}\left (t\right )\left (1-t\right )+\mathcal{U}\left (t-1\right )\left (t-1\right )\]

OpenStudy (anonymous):

Let's check it.

OpenStudy (anonymous):

\[f(t)=\begin{cases} -t+1 & \text{ if } 0\leq t< 1 \\ 0 & \text{ if } t\geq 1 \end{cases}\]

OpenStudy (anonymous):

\[\mathcal{L}\left \{ f(t) \right \}=\int_{0}^{1}e^{-st}(-t+1)dt+\int_{1}^{\infty}e^{-st}(0)dt\]

OpenStudy (anonymous):

\[\mathcal{L}\left \{ f(t) \right \}=-\int_{0}^{1}e^{-st}tdt+\int_{0}^{1}e^{-st}dt+\int_{1}^{\infty}e^{-st}(0)dt\]

OpenStudy (across):

That's correct.